OpenStudy (anonymous):

cant solve this, can anyone help 2015 = x^2(5-x)^2

OpenStudy (anonymous):

\[x^2+(5-x)^2 = 2015\]

OpenStudy (anonymous):

Expand the left side, then move all values to the left side to make a quadratic equation.

OpenStudy (anonymous):

\[2x^2-10x+25 = 2015\]

OpenStudy (anonymous):

\[(a + b)^2 = a^2 + b^2 + 2*a*b\] Use this and make it small. After that use this \[ax^2+bx + c = 0 \\ x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

x^2+25-10x+x^2 = 2015 2(x^2)-10x+25-2015 = 0 2(x^2)-10x+1990 = 0 x^2-5x+995=0

OpenStudy (anonymous):

isint it -995?

OpenStudy (anonymous):

oops yeah sorry -995

OpenStudy (anonymous):

okay got it thank you very much!!